OpenStudy (anonymous):

Find the distance for the following points A ( -2, -3) B (-7, -8). If necessary round your answer to the nearest tenth. 1) 11 2) 7.1 3) 7 4) -7.1

OpenStudy (bahrom7893):

distance = sqrt ( (x1-x2)^2 + (y1-y2)^2 ) = = sqrt ( (-7 - [-2])^2 + ( -8 - [-3])^2 )

OpenStudy (anonymous):

You know immediately that you can knock out number 4 because a distance will never been negative :P Then do as bahrom said^^

OpenStudy (bahrom7893):

= sqrt ( (-7+2)^2 + (-8+3)^2 ) = sqrt ( (-5)^2 + (-5)^2 ) = sqrt ( 25 + 25 ) = sqrt (50)

OpenStudy (bahrom7893):

which is approximately 7.1